OpenStudy (anonymous):

if a^2 + a + 1=0 then a^9 is?

OpenStudy (freckles):

You could use the quadratic formula to find a then raise it to the 9th power.

OpenStudy (freckles):

I'm not sure if there should be a trick here though or not...

OpenStudy (freckles):

I will try to think of a trick.

OpenStudy (freckles):

I suppose this is a trig question actually.

OpenStudy (aum):

You can complete the square on the left, solve for a and then raise it to the 9th power.

OpenStudy (freckles):

You could Solve a^2+a+1=0 Then a=y+zi=e^{i theta} theta is easy to find in this case

OpenStudy (freckles):

y and z are real numbers of course

OpenStudy (freckles):

so you would have a=e^{i (theta} Raise both sides the the 9th power

OpenStudy (freckles):

that would be easier than doing a^9 algebraically

OpenStudy (zarkon):

\[a^2+a+1=0\] so \[a^9+a^8+a^7=0\Rightarrow a^9=-(a^8+a^7)\] but \[a^6\times(a^2+a+1=0)\Rightarrow a^8+a^7=-a^6\] thus \[a^9=-(a^8+a^7)=-(-a^6)=a^6\] ...keep doing this.

OpenStudy (anonymous):

it was asked in entrance exam . and had 1 mins to solve

OpenStudy (freckles):

Well I suggested a trig way. Do you know trig?

OpenStudy (zarkon):

it only takes about a min ;)

OpenStudy (freckles):

That is an interesting way zarkon

OpenStudy (zarkon):

there is a pattern ... \[a^9=a^6=a^3=a^0=1\]

OpenStudy (freckles):

a^9=-1

OpenStudy (zarkon):

no...it is 1

OpenStudy (zarkon):

to verify I just solved it on my calculator and raised the answer to the 9th power...got 1

OpenStudy (freckles):

oops I had 1/2+- sqrt(3)/2

OpenStudy (freckles):

forgot to put - in front of the 1/2

OpenStudy (freckles):

\[a=\frac{-1}{2} \pm \frac{\sqrt{3}}{2} \\ a=e^{\frac{2\pi}{3} i} \\ a^9=e^{6\pi i}\]

OpenStudy (freckles):

=1

OpenStudy (freckles):

sorry my bad stupid -

OpenStudy (zarkon):

that seems like a lot of work for this problem :)

OpenStudy (freckles):

That isn't a lot of work if you know the unit circle

OpenStudy (freckles):

But that is my opinion.

OpenStudy (anonymous):

thanks all :)

hartnn (hartnn):

a^2 + a + 1=0 (a-1) (a^2+a+1) =0 a^3 -1 = 0 a^3 =1 a=1 a^9 =1 took about 4 secs

OpenStudy (zarkon):

you can't say \(a=1\) (because it is not...that zero was introduced when you multiplied by a-1) but you can go from \(a^3=1\) to \((a^3)^3=(1)^3\Rightarrow a^9=1\)

hartnn (hartnn):

right! sorry, my bad

OpenStudy (zarkon):

nice idea though.